OpenStudy (anonymous):

Write the standard form of the equation of the line that passes through the point (4,1) and is parallel to the line 3x + 2y = 5

OpenStudy (texaschic101):

first you need to put 3x + 2y = 5 into y = mx + b form. That is how we will find the slope. Can you do this ?

OpenStudy (anonymous):

uhm.. can you give me the answer and explain it..

OpenStudy (texaschic101):

3x + 2y = 5 (subtract 3x from both sides) 2y = -3x + 5 (now divide by 2) y = -3/2x + 5/2 In the y = mx + b form, the m stands for the slope. Therefore,the slope in this equation is -3/2. If you are looking for a parallel line, the slope is the same as the slope in your equation. Now we will use this formula : y - y1 = m(x - x1) slope(m) = -3/2 (4,1) x1 = 4 an y1 = 1 now lets sub... y - 1 = -3/2(x - 4) -- distribute through the parenthesis y - 1 = -3/2x + 6 -- add 1 to both sides y = -3/2x + 6 + 1 -- combine like terms y = -3/2x + 7 <-- slope intercept form Now we need to rearrange this equation and put it in standard form y = -3/2x + 7 -- add 3/2x to both sides 3/2x + y = 7 -- multiply the equation by 2 to get rid of fractions 3x + 2y = 14 <-- standard form any questions ?

OpenStudy (anonymous):

let me go though this. :) thanks in advance sir

OpenStudy (texaschic101):

not sir...I am a lady...lol ook through it and if you have any questions, tag me :)

OpenStudy (texaschic101):

typo * look

OpenStudy (anonymous):

oo.. my mistake,.. im sorry madam..

OpenStudy (anonymous):

i assume that this 3x + 2y = 5 (subtract 3x from both sides) will look like this. -3x+3x+2y=5

OpenStudy (anonymous):

-3x+3x+2y=-3x+5

OpenStudy (texaschic101):

yes....I just didn't put EVERY step in.

OpenStudy (anonymous):

3/2x + y = 7 -- multiply the equation by 2 to get rid of fractions how to multiply this?

OpenStudy (texaschic101):

2(3/2)x + 2(y) = (2)7 6/2x + 2y = 14 3x + 2y = 14 you just multiply by the LCD to get rid of the fractions

OpenStudy (texaschic101):

you understand it ?

OpenStudy (anonymous):

yes :DD

OpenStudy (anonymous):

thank you very much madam. ^^

OpenStudy (texaschic101):

lol...no problem :)

OpenStudy (anonymous):

can i ask a few more questions?